### v2

# Ocean-E2E: Hybrid Physics-Based and Data-Driven Global Forecasting of Extreme Marine Heatwaves with End-to-End Neural Assimilation

Abstract
This work focuses on the end-to-end forecast of global extreme marine heatwaves (MHWs), which are unusually warm sea surface temperature events with profound impacts on marine ecosystems. Accurate prediction of extreme MHWs has significant scientific and financial worth. However, existing methods still have certain limitations, especially in the most extreme MHWs. In this study, to address these issues, based on the physical nature of MHWs, we created a novel hybrid data-driven and numerical MHWs forecast framework Ocean-E2E, which is capable of 40-day accurate MHW forecasting with end-to-end data assimilation. Our framework significantly improves the forecast ability of extreme MHWs by explicitly modeling the effect of oceanic mesoscale advection and air-sea interaction based on a differentiable dynamic kernel. Furthermore, Ocean-E2E is capable of end-to-end MHWs forecast and regional high-resolution prediction using neural data assimilation approaches, allowing our framework to operate completely independently of numerical models while demonstrating high assimilation stability and accuracy, outperforming the current state-of-the-art ocean numerical forecasting-assimilation models. Experimental results show that the proposed framework performs excellently on global-to-regional scales and short-to-long-term forecasts, especially in those most extreme MHWs. Overall, our model provides a framework for forecasting and understanding MHWs and other climate extremes. Our codes are available at https://github.com/ChiyodaMomo01/Ocean-E2E.
1 Introduction
Marine heatwaves (MHWs) are abnormally warm seawater events that significantly damage marine ecosystems oliver2021marine ; pearce2011marine . In other words, MHWs are extreme Sea Surface Temperature Anomaly (SSTA) events. MHWs, particularly extreme ones, can cause coral bleaching hughes2017global ; hughes2018spatial and widespread mortality of marine organisms garrabou2009mass ; thomson2015extreme . As a result, precise forecasting of extreme MHWs has significant scientific and economic implications. For example, synoptic scale MHWs forecasting can help sea food production and management planning, such as feed cycles, at 1-7 day timescales, while subseasonal to seasonal forecasting can further support proactive decision-making for the blue economy hobday2016seasonal ; malick2020environmentally ; mills2017forecasting ; payne2022skilful . In this study, we will mainly focus on the subseasonal-to-seasonal forecast (i.e., 1-40 day).
Traditional MHW forecasting predominantly relies on numerical models solving oceanic primitive equations, which generally bifurcate into two categories: seasonal forecasts characterizing large-scale patterns of MHW onset and intensification jacox2022global ; brodie2023ecological , and sub-seasonal predictions resolving finer-scale variability with reduced lead times benthuysen2021subseasonal ; yu2024assessing . While these physics-based approaches achieve operational utility, they demand substantial computational resources xiong2023ai ; wang2024xihe and exhibit limited skill in predicting extreme MHW events jacox2022global . Recent advances in deep learning have demonstrated potential for efficient global ocean forecasting bi2022pangu ; chen2023fengwu ; kurth2023fourcastnet , with emerging applications to MHW prediction lin2023rossby ; sun2023artificial . However, existing implementations face dual limitations: 1) insufficient encoding of MHW-specific physics—particularly air-sea coupling mechanisms driving extreme events and mesoscale eddy dynamics governing local intensification—resulting in non-physical smoothing of SST gradients during autoregressive prediction giamalaki2022assessing ; 2) dependence on numerical model initialization that prevent operational end-to-end forecasting from raw observational inputs xiong2023ai . These architectural constraints—shared across mainstream deep learning paradigms—fundamentally limit their capacity to preserve extreme SST anomalies and long-term physical consistency, as error accumulation from neural networks and initialization biases during long-term rollouts wang2024xihe .
By addressing the two shortcomings of the data-driven methods mentioned above, in this study, we designed Ocean-E2E, a global data-driven framework that significantly improves the forecasts of extreme MHWs inspired by numerical models. The contribution of this paper can be summarized as follows: (1) Global MHWs Forecasting Framework. We propose a hybrid physics-based data-driven method that supports both global scale and regional high-resolution forecasting, achieving high-accuracy results for extreme MHWs forecasts. (2) Global MHWs Neural Assmilation Framework. We designed an end-to-end MHWs neural system that, when combined with the forecasting model, can directly obtain the complete initial field and forecasting results of MHWs from observation data. (3) Long-Term Forecast. By combining phyiscal laws into our framework, our model has achieved state-of-the-art results in long-term (subseasonal-to-seasonal) forecasting with higher physical consistency.
The method most similar to us is ClimODE verma2024climode . Similarly, it also uses advection partial differential equations to design the forecasting framework. However, our work has many differences and improvements compared to ClimODE: (1) Differences in theoretical framework. From the perspective of the theoretical framework, in their work, the ‘speed’ of the flow field is essentially an abstract mathematical definition, whose core function is to maintain the conservation of passive tracers. In our work, the flow field speed strictly corresponds to the realistic ocean flow field and models the additional transport effect of the flow field on passive tracers at sub-grid scales through mesoscale ‘bolus’ velocity. (2) Differences in model design. ClimODE uses an emission model to simulate the unclosed and uncertain portions of the real system, while our model further relates this part to the ocean-atmosphere interaction and is designed based on physical theory to form the structure of the source term. (3) Differences in application. ClimODE requires solving the initial velocity field as the initial condition for the model, which involves solving a physics-informed neural network (PINNs) for each training sample. This makes the ClimODE framework difficult to scale up. What’s more, our experiments (Section 4.3) show that our method achieves greater stability and accuracy in MHWs forecasting compared to ClimODE.
Figure 1: Our proposed Ocean-E2E framwork. a) Overview of our framework. b) Details of our hybrid physics and data-driven forecast model.
2 Problem Setup
In this study, the problem of the end-to-end MHWs forecast is essentially forecasting SSTA, which can be separated into two parts. First, estimating the current state of SSTA $T_{0}\in\mathbb{R}^{H\times W}$ (a global two-dimensional field, where $H$ and $W$ correspond to latitude and longitude) based on the observation $T_{0}^{obs}$ and the background field $T_{0}^{bg}$. Second, predicting the future state of mixed layer SSTA $T_{t}$ based on the current state $T_{0}$ under atmospheric forcing $A_{0:t}\in\mathbb{R}^{T\times H\times W}$. Here the underscore $0:t$ denotes the atmosphere state from time 00 to $t$. Formally speaking, we are going to model the conditional distribution $p(T_{t}|T_{0}^{obs},T_{0}^{bg},A_{0:t},)$ in two stages:
$p(T_{t}|T_{0}^{obs},T_{0}^{bg},A_{0:t})=\underbrace{p^{f}(T_{t}|T_{0},A_{0:t})}_{\text{Forecast model}}\cdot\underbrace{p^{a}(T_{0}|T_{0}^{obs},T_{0}^{bg})}_{\text{Assimilation model}}$ (1)
where $p^{f}$ and $p^{a}$ denote forecast and assimilation process, respectively. In our end-to-end framework, we model these two processes through two specifically designed neural network $\phi^{f}_{\theta}$ and $\phi^{a}_{\theta}$:
$T_{t}=\phi^{f}_{\theta}(T_{0},A_{0:t})$ (2)
$T_{0}=\phi^{a}_{\theta}(T_{0}^{obs},T_{0}^{bg})$ (3)
In summary, our end-to-end framework aims to forecast the furture state of MHWs $T_{t}$ based on the sparse observation $T_{0}^{obs}$ and the initial background field $T_{0}^{bg}$ under the forcing of atmosphere boundary conditions $A_{0:t}$. The parameters of these two neural networks $\theta_{f},\theta_{a}$ are optimized separately by minizing the mean square error (MSE) between the model’s output and the groundtruth:
$\theta_{f}={\rm argmin}\{\frac{1}{N}\sum_{i=0}^{N}(T_{t}^{(i)}-\phi^{f}_{\theta_{f}}(T_{0}^{(i)},A_{0:t}^{(i)}))^{2}\}$ (4)
$\theta_{a}={\rm argmin}\{\frac{1}{N}\sum_{i=0}^{N}(T_{t}^{(i)}-\phi^{a}_{\theta_{a}}(T_{0}^{(i)},A_{0:t}^{(i)}))^{2}\}$ (5)
3 Methods
3.1 Motivation and Framework Overview
Our framework is summarized in Figure 1. As established in the problem formulation, forecasting MHWs fundamentally reduces to predicting extreme sea surface temperature anomalies (SSTA). To effectively embed physical principles into our modeling framework, we must rigorously account for the governing laws of SSTA evolution. We approach this challenge through a generalized geophysical fluid dynamics perspective, wherein seawater temperature constitutes a passive tracer $C$ advected by oceanic flow fields. As shown in Figure 2, the spatiotemporal evolution of $C$ is governed by two primary mechanisms: 1) advective transport by background currents that redistribute thermal energy, 2) mixing and external sink/source encompassing convective mixing and external forcing through surface boundary interactions. Formally speaking, the governing equation of $C$ can be written as:
$\frac{\partial C}{\partial t}+{\bf u}\nabla C+w\frac{\partial C}{\partial z}=\dot{\rm S}+\nabla\cdot(\kappa\nabla C)$ (6)
Where $t$ is time, $({\bf u},w)$ are horizontal and vertical flow velocity, $\nabla$ is the horizontal derivation, $\dot{\rm S}$ is the sink/source term and $\kappa$ is a $2\times 2$ matrix corresponding to horizontal anistropic diffusion coeffiecient smith2004anisotropic . By integrating the above equation, the state of $C$ at time $t$ can be expressed in the form of the initial condition $C_{0}$:
$C_{t}=\int_{0}^{t}-{\bf u}\nabla C+(-w\frac{\partial C}{\partial z}+\dot{\rm S}+\nabla\cdot(\kappa\nabla C))dt+C_{0}$ (7)
Figure 2: Governing Mechanisms of MHWs
Where $-{\bf u}\nabla C$ represents the advective transport while $-w\frac{\partial C}{\partial z}+\dot{\rm S}+\nabla\cdot(\kappa\nabla C)$ corresponds to mixing and external sink/source. Next, we will combine physical theory and deep learning methods to simulate these two mechanisms respectively.
Advective transport. The key factor of advective transport is the ocean surface current ${\bf u}$, where the flow field at the smaller scale (mesoscale) plays an important role in the distribution of MHWs and other oceanic passive tracers. However, in most models and data, due to limited resolution, the model cannot distinguish these small-scale processes, so the advection processes caused by the flow field below the model’s finest grid (i.e., subgrid processes):
$\overline{{\bf u}\nabla C}=\overline{{\bf u}}\overline{\nabla C}+\overline{{\bf u}^{\prime}{\nabla C}^{\prime}}$ (8)
are not fully simulated. Here the overline $\overline{\cdot}$ represents the downsampling operator. In other words, the variables with overline denote the low-resolution data. In the traditional numerical model, Gent and McWilliams gent1990isopycnal found that this subgrid process can be fitted by ‘bolus velocity’ (also called GM90 parameterization):
$\overline{{\bf u}^{\prime}{\nabla C}^{\prime}}\sim{\bf u}_{gm}\overline{\nabla C}$ (9)
According to the GM90 parameterization, these bolus velocity can be expressed asgent1990isopycnal :
${\bf u}_{gm}=(\kappa\nabla\rho/\rho_{z})_{z}$ (10)
where $\rho=\rho(T)$ is the density. It should be noticed that this expression requires subsurface ocean data, which is often extremely difficult to obtain. What’s more, the choice of coefficient $\kappa$ requires empirical formula which is always inaccurate. Alternatively, we seek for a neural network ${\bf u}_{\theta}({\bf u},C)={\bf u}_{gm}$ to approximate the bolus velocity and the subgrid advection effect. In summary, the advection term can be expressed as:
${\bf u}\nabla C=({\bf u}_{\theta}({\bf u},C)+{\bf u})\nabla C$ (11)
Mixing and external sink/source. This part consists of three components: verical mixing, external source, and dissipation. Vertical mixing and dissipation depend more on the instrinic variations of the ocean while the external source $\dot{\rm S}$ is strongly related to the air-sea interaction along the ocean’s surface. To be specific, for SSTA, the source term consists of four parts:
$\dot{\rm S}\propto Q_{{\rm net}}^{\prime}=Q_{{\rm SW}}^{\prime}+Q_{{\rm LW}}^{\prime}+Q_{{\rm sh}}^{\prime}+Q_{{\rm lh}}^{\prime}$ (12)
where $Q_{{\rm SW}},Q_{{\rm LW}},Q_{{\rm sens}},Q_{{\rm lat}}$ is the shortwave radiation, longwave thermal radiation, sensible heat flux, and latent heat flux respectively. Of the four air-sea heat fluxes, the sensible and latent flux $Q_{{\rm sens}}^{\prime}+Q_{{\rm lat}}^{\prime}$ determines the majority of the SSTA’s variations holbrook2019global . According to the bulk formula, these two terms are significantly related to surface wind speed $U_{10},V_{10}$, near-surface temperature of the atmosphere $T_{2m}$, and surface specific humidity $q_{a}$. Based on the observation above, we assume that mixing and external sink/source can be well approximated by the surface variables of the atmosphere ${\bf A}=(U_{10},V_{10},T_{2m},q_{a})$ and $C$ itself. Specifically,
$\dot{\rm S}_{\theta}({\bf A},C)=-w\frac{\partial C}{\partial z}+\dot{\rm S}+\nabla\cdot(\kappa\nabla C)$ (13)
In summary, the disscusions above illustrate that the governing processes of SSTA can be efficiently modelled by neural networks. In the following part, we will demonstrate how to incorporate this idea into a hybrid physics-AI model. Specifically, for the bolus velocity approximation ${\bf u}_{\theta}$, we employ a U-Net like architecture. The core of this network, particularly in its deeper layers or bottleneck, utilizes GABlocks (composed of GASubBlock units which integrate SpatialAttention and Mlp modules) to capture multi-scale spatial dependencies from the input ocean current ${\bf u}$ and SSTA $C$. This structure allows the model to learn complex, non-local relationships crucial for representing subgrid advection effects. For the mixing and external sink/source term $\dot{\rm S}_{\theta}$, we utilize a network primarily built upon a stack of GABlocks. This network takes atmospheric variables ${\bf A}$ and the current SSTA $C$ as input, enabling it to effectively model the intricate air-sea interactions and other relevant source/sink processes. The GABlock, with its SpatialAttention mechanism (itself using depth-wise and dilated convolutions as seen in AttentionModule) and Mlp (Multilayer Perceptron with depth-wise convolution, DWConv), is designed to efficiently process and integrate these diverse spatio-temporal inputs for both ${\bf u}_{\theta}$ and $\dot{\rm S}_{\theta}$. The details structure of ${\bf u}_{\theta}$ and $\dot{\rm S}_{\theta}$ can be found in Appendix D.
3.2 Hybrid physics-AI forecast model
As shown in the section above, the furture state of $C$ at time $t$ can be modelled as:
$C_{t}=\int_{0}^{t}-({\bf u}(t)+{\bf u}_{\theta}({\bf u}(t),C)\nabla C+\dot{\rm S}_{\theta}({\bf A}(t),C)dt+C_{0}$ (14)
This indicates that if we are given the values of the ocean currents and atmospheric boundary conditions between the time interval $[0,t]$, we can predict $C_{t}$ in a hybrid physics-AI way.
In our study, the ocean current is obtained from surface geostrophic current inferred from satellite observations. The temporal evolution of this geostrophic current is mainly driven by the ocean instrinic mesoscale dynamics, which can be rigorously described by the 1.5-layer nonlinear RGO model capotondi2001rossby . In other words, the evolution of ${\bf u}$ can be modelled based on an auto-regressive manner:
$\frac{\partial{\bf u}}{\partial t}=\mathcal{M}_{\theta}({\bf u})$ (15)
What’s more, the atmosphere variables between time interval $[0,t]$ is also acquired through a large AI-driven weather forecast model gao2025oneforecast :
$\frac{\partial{\bf A}}{\partial t}=\mathcal{N}_{\theta}({\bf A})$ (16)
Here $\mathcal{M}_{\theta},\mathcal{N}_{\theta}$ are two pretrained neural networks. Details of these neural networks can be found in Appendix D. In all, the forecast model can be expressed as:
[…]
In the training stage, we first pretrain these two neural networks $\mathcal{M}_{\theta},\mathcal{N_{\theta}}$. Then the parameters of these networks are frozen and we utilize the forecast model to optimize the parameters of ${\bf u}_{\theta},\dot{\rm S}_{\theta}$ based on the MSE loss, which can be written as:
$\mathcal{L}=||\hat{C_{t}}-\int_{0}^{t}-({\bf u}+{\bf u}_{\theta}({\bf u},C))\nabla C+\dot{\rm S}_{\theta}(A,C)dt||^{2}$ (18)
Where $\hat{C_{t}}$ is the groundtruth data of $C_{t}$. In practice, by separating the time interval $[0,t]$ into $N$ subintervals $[i\Delta t,(i+1)\Delta t],1\leq i\leq N-1$, equation (18) is approximated via the second-order Runge-Kutta method:
$\tiny C_{(i+1)\Delta t}^{\star}$ $=C_{i\Delta t}+\left[-\left({\bf u}_{i\Delta t}+{\bf u}_{\theta}({\bf u}_{i\Delta t},C_{i\Delta t})\right)\nabla C_{i\Delta t}+\dot{\rm S}_{\theta}(A_{i\Delta t:(i+1)\Delta t},C_{i\Delta t})\right]\Delta t$ (19)
$C_{(i+1)\Delta t}$ $=C_{i\Delta t}+\frac{1}{2}\Bigg{\{}\left[-\left({\bf u}_{i\Delta t}+{\bf u}_{\theta}({\bf u}_{i\Delta t},C_{i\Delta t})\right)\nabla C_{i\Delta t}+\dot{\rm S}_{\theta}(A_{i\Delta t:(i+1)\Delta t},C_{i\Delta t})\right]\Delta t$
$\quad+\left[-\left({\bf u}_{(i+1)\Delta t}+{\bf u}_{\theta}({\bf u}_{(i+1)\Delta t},C_{(i+1)\Delta t}^{\star})\right)\nabla C_{(i+1)\Delta t}^{\star}+\dot{\rm S}_{\theta}(A_{(i+1)\Delta t:(i+2)\Delta t},C_{(i+1)\Delta t}^{\star})\right]\Bigg{\}}$ (20)
3.3 End-to-end Neural Data Assimilation
Under realistic operational forecasting scenarios, the full initial ocean state $T_{0}$ (analysis field) cannot be directly observed. This necessitates a data assimilation (DA) process that optimally combines sparse observational data $T_{0}^{obs}$ with prior background estimates $T_{0}^{bg}$ from numerical models. Traditional DA methodologies require large computational resources, which imposes an inherent trade off between solution accuracy and operational feasibility boudier2023data .
Our framework addresses these limitations by establishing a direct mapping from observational inputs $T_{0}^{obs}$ to assimilated states $T_{0}$ through deep learning. Instead of solving complex Bayesian inverse problems, we train neural networks to optimally combine sparse observations with background fields in one step. Specifically, we begin with the SSTA state at time $-t$ ($T_{-t}$), defining $[-t,0]$ as the assimilation window. We generate background fields ($T_{0}^{bg}$) by adding structured noise to $T_{-t}$:
$T_{0}^{bg}=\phi^{f}_{\theta}(T_{-t}+\varepsilon,A_{0:t})$ (21)
where $\varepsilon$ denotes Perlin noises. The final analysis field is then produced by merging observations with background data:
$T_{0}=\phi^{a}_{\theta}(T_{0}^{obs},T_{0}^{bg})$ (22)
Where $\phi^{a}_{\theta}$ is an edge-reparameterization and attention-guided network with Kirsch-guided reparameterization, which is capable of reconstructing SSTA fields from sparse observations liu2024real . The details of $\phi^{a}_{\theta}$ can be found in Appendix D.
4 Experiment
4.1 Benchmarks and Baselines
We conduct the experiments on GLORYS12V1 reanalysis data lellouche2018copernicus , which provides daily mean data of sea surface temperature (SST) covering latitudes between -80° and 90° spanning between 1993 and 2021. The seasonal cycle has been removed from the original data in order to get the MHWs (SSTA) field shu2025advanced . The subset we use includes years from 1993 to 2021, which is 1993-2018 for training, 2019 for validing, and 2020-2021 for testing. The surface velocity of the ocean is obtain from satellie observation pujol2016duacs . The atmospheric variables are obtained from ECWMF Reanalysis v5 (ERA5) hersbach2020era5 . As shown in section 3.1, based on the physical governing process, we select three surface atmosphere variables to represent atmsopheric forcing $A_{0:t}$: 10m wind components ($U_{{\rm 10}}$, $V_{{\rm 10}}$) and 2m temperature ($T_{{\rm 2m}}$). More details, including observational data, can be found in Appendix B.
Table 1: In global ocean simulation task, we compare the performance of our Ocean-E2E with 4 baselines. The average results for global SSTA of RMSE and CSI are recorded. A small RMSE ($\downarrow$) and a higher CSI ($\uparrow$) indicate better performance. The best results are in bold, and the second best are with underline. It should be noticed that some of the baseline models are not stable in this task and generate unrealisitic forecast results. As a result, RMSE exceeds 10 K is replaced with --.
Model Metric
20-day 40-day 50-day 60-day
RMSE CSI RMSE CSI RMSE CSI RMSE CSI
FourCastNet pathak2022FourCastNet 0.6836 0.2709 0.1755 0.1537 0.1423
CirT liu2025cirt 1.3496 0.0905 1.9249 0.0810 1.9932 0.0770 2.0321 0.0746
WenHai cui2025forecasting 0.5435 0.4202 0.7006 0.2805 0.7633 0.2447 0.8139 0.2209
ClimODE verma2024climode 0.7221 0.3263 0.8555 0.2091 0.8997 0.1896 0.1754
Ocean-E2E (ours) 0.5659 0.4285 0.6596 0.3230 0.6911 0.2874 0.7194 0.2580
Ocean-E2E (Promotion) 2.0% 5.8% 15.1% 9.5% 17.4% 11.6% 16.8%
Before introducing our experiments, there are two key concepts that should be clarified: ocean simulation and end-to-end ocean forecast. As pointed out in previous studiesshu2025advanced ; cui2025forecasting , the ocean simulation utilizes the current state of SSTA $T_{0}$ to forecast future state $T_{t}$ under the realistic atmosphere forcing ${\bf A}_{0:t}$. In contrast, the end-to-end forecast task includes assimilating the initial analysis field of SSTA $T_{0}$ and forecasting only based on the current atmosphere state ${\bf A}_{0}$.
We compare the simulation performance of Ocean-E2E with 3 data-driven models (Fourcastnetkurth2023fourcastnet , CirTliu2025cirt , and Wenhaicui2025forecasting ). Fourcastnet is an atmosphere backbone model with robust performance, CirT is a SOTA subseasonal-to-seasonal atmosphere prediction model, and Wenhai is the superior data-driven ocean model that outperforms traditional numerical models and data-driven models.
We also compare the end-to-end operational forecast ability against traditional numerical operational forecast system: ECWMF sub-seasonal to seasonal prediction (S2S). S2S is current superior subseasonal-to-seasonal weather forecast model that provides the forecast of atmosphere and ocean surface physics at a lead time of 40 days. The original resolution is 1.5° and we upsample the data to 1/2° for comparison.
4.2 Evalution Metric
We utilize two metrics, RMSE (Root Mean Square Error) and CSI (Critical Success Index), to evaluate the forecast performance. The RMSE represents the overall performance of our framework’s forecast, while CSI mainly focuses on the most extreme MHWs events. What’s more, we utilize assimilation bias (BIAS) to assess the systematic bias induced by our neural assimilation system. More details can be found in Appendix E.
Figure 3: Snapshots of our framework and other baseline simulations at 60-day lead time. The simulations are initialized on August 7th, 2020.
Table 2: In operational global ocean forecast task, we compare the performance of our Ocean-E2E with S2S. The average results for global SSTA of RMSE are recorded. A small RMSE ($\downarrow$) indicates better performance. The best results are in bold, and the second best are with underline.
Model Metric (RMSE)
Time Interval
10-day 20-day 30-day 40-day
S2S (2020) pathak2022FourCastNet 0.8140 0.8870 0.9514 0.9965
Ocean-E2E (2020) liu2025cirt 0.5750 0.6047 0.7514 0.8747
Ocean-E2E (Promotion) 29.4 % 31.8 % 21.02 % 12.2 %
S2S (2021) pathak2022FourCastNet 0.8261 0.8897 0.9450 0.9895
Ocean-E2E (2021) liu2025cirt 0.6414 0.7447 0.8147 0.8729
Ocean-E2E (Promotion) 22.4 % 16.3 % 13.7 % 11.7 %
Figure 4: Results of our end-to-end assimilation framework. a) Snapshots of our analysis field and the groundtruth after 300 days of the initialization of our assimilation in 2020 b) Temporal evolution of RMSE and BIAS during the assimilation period.
Figure 5: An illustration of the regional high-resolution simulation.
4.3 Evalution of ocean simulation performance
[…]
4.4 Evaluation of the end-to-end neural assimilation and forecast performance
We conducted two sets of assimilation experiments using analysis fields from January 1, 2020 and January 1, 2021 as initial conditions, respectively. Both experiments employed a 6-day assimilation window to assimilate SSTA fields over a full year. As shown in Figure 4, the RMSE and BIAS evolution reveal that assimilation errors rapidly increase and stabilize during the assimilation process, demonstrating the stability of our assimilation framework. Notably, a persistent negative BIAS exists in our framework, indicating that our assimilation model systematically underestimates the intensity of marine heatwaves compared to GLORYS12V1 reanalysis data, as also evidenced in Figure 4a.
[…]
4.5 Regional high-resolution simulation
As discussed in Section 3.1, our framework builds upon the GM90 theoretical framework, which is applicable to all eddy-permitting datasets (horizontal resolution: 10–100 km). To validate its performance, we implement a regional high-resolution simulation ($1/4^{\circ}$) of the Pacific Ocean. Figure 5 depicts the evolution of sea surface temperature anomaly (SSTA) during a 20-day simulation, governed by the equation:
$C_{t}-C_{0}=\int_{0}^{t}-\left(\mathbf{u}(t)+\mathbf{u}_{\theta}(\mathbf{u}(t),C)\nabla C+\dot{S}_{\theta}(\mathbf{A}(t),C)\right)dt,$ (23)
where $C_{t}$ and $C_{0}$ represent the SSTA at time $t$ and initial state, respectively. The results demonstrate that Ocean-E2E effectively resolves multi-scale physical processes: large-scale SSTA patterns align with air-sea interactions, while small-scale variations correspond to local advection effects quantified by $\int_{0}^{t}-\left(\mathbf{u}(t)+\mathbf{u}_{\theta}(\mathbf{u}(t),C)\right)\nabla C\,dt$. By explicitly modeling these components, the framework achieves rigorous reconstruction of $C_{t}-C_{0}$, confirming its capability to disentangle scale-dependent dynamics in MHWs bian2024scale .
Table 3: CSI of our ablation studies on model design, the best results are in bold.
Variants 20 day 60 day
Ocean-E2E w/o ADV 0.3884 0.1658
Ocean-E2E w/o S 0.3207 0.1504
Ocean-E2E 0.4285 0.2580
4.6 Ablation studies
To validate the effectiveness of the proposed framework, we perform an ablation analysis by sequentially removing two core components: the advection term $({\bf u}_{\theta}({\bf u},C)+{\bf u})\nabla C$ (denoted as ADV) and the source term $\dot{S}_{\theta}(\mathbf{A},C)$ (denoted as S). As demonstrated in Table 3, the Critical Success Index (CSI) exhibits a marked reduction when either component is omitted, confirming their essential roles in enhancing simulation accuracy, particularly for extreme marine heatwaves (MHWs). Notably, the source term contributes most significantly to forecasting performance, underscoring the critical influence of air-sea interaction dynamics on MHW evolution.
5 Conclusion
In this study, we present Ocean-E2E, a novel hybrid framework integrating data-driven and numerical approaches for end-to-end marine heatwave (MHW) forecasting. By explicitly resolving oceanic mesoscale advection and air-sea interactions through a differentiable dynamic kernel, the framework achieves stable 40-day global-to-regional MHW predictions. Notably, in 60-day global ocean simulations, Ocean-E2E demonstrated a significant improvement, reducing the RMSE by an average of 11.6% and increasing the CSI for extreme MHWs by an average of 9.2% compared to leading baseline models. Furthermore, in end-to-end operational forecast tasks, our framework outperformed the state-of-the-art S2S system, achieving an average RMSE reduction of approximately 10% over a 40-day forecast horizon using its own assimilated initial fields. Neural data assimilation further enables high-resolution forecasting independent of traditional numerical systems, while maintaining superior accuracy and stability, as evidenced by the rapid stabilization of assimilation errors. Experimental results comprehensively demonstrate its robustness across spatiotemporal scales, offering a promising tool for advancing MHW prediction and understanding climate extremes.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (42125503, 42430602).
Appendix A Algorithm
We summarize the overall framework of Ocean-E2E in Algorithm 1.
Algorithm 1 Ocean-E2E Framework for Global MHWs Forecast
0:  Initial SSTA observation $T_{t}^{obs}$, analysis field $T_{t-1}^{a}$ (at time $t-1$), current state of atmospheric forcing $\mathbf{A}_{t}$ and ocean current velocity $\mathbf{O}_{t}$
0:  Current step SSTA analysis field $T_{t}^{a}$, Next step SSTA prediction $T_{t+1}$.
1:  Initialize Ocean-E2E
2:  repeat
3:     Data Assimilation
4:     Perturb the initial SSTA field $T_{t-1}^{a}$ with $N=10$ Perlin noise $T_{t-1}^{a(i)}=T_{t-1}^{a}+\varepsilon^{(i)},i=1,2,...,N$.
5:     Feeding the perturbed ensembles into our trained forecast model to generate forecast ensemble: $T_{t}^{bg(i)}=\phi^{f}_{\theta}(T_{t-1}^{a(i)},\mathbf{A}_{t-1:t})$, which are also called background fields at time $t$.
6:     Using the neural assimilation network $\phi^{a}_{\theta}$ to produce the analysis field at time $t$: $T_{t}^{a}=\phi^{a}_{\theta}(T_{t}^{bg(i)},T_{t}^{obs})$
7:     Hybrid physics data-driven forecast
8:     Calculate the future atmospheric forcing and ocean current velocity: $\mathbf{A}_{t:t+1}=\mathcal{M}_{\theta}(\mathbf{A}_{t}),\mathbf{O}_{t:t+1}=\mathcal{M}_{\theta}(\mathbf{O}_{t})$.
9:     Taking the current analysis field $T_{t}^{a}$, and integrating the equation (17) with $\mathbf{A}_{t:t+1}$ and $\mathbf{O}_{t:t+1}$ to get the forecast field $T_{t+1}$
10:  until converged
11:  return $Ocean-E2E$
Appendix B Data Details
B.1 Dataset
In this section, we will introduce the dataset used in this study in detail.
Data used to train $\mathcal{M}_{\theta}$. As mentioned in Section 3.2, we need to pretrain an neural network $\mathcal{M}_{\theta}$ to model the evolution of the atmosphere forcing $\mathbf{A}$. In this study, we utilize one of the state-of-the-art weather forecast model: OneForecast [17]. The data used to train OneForecast is obtained from weatherbench2 [48] benchmark, which is a subset of ERA5 reanalysis data [24]. Details can be found in Table 4:
Table 4: The data used to train $\mathcal{M}_{\theta}$.
Variable
Name
Layers
Spatial
Resolution
Dt
Lat-Lon
Range
Time
Geopotential (Z) 13 1.5° 6h Global 1959$\sim$2021
Specific Humidity (Q) 13 1.5° 6h Global 1959$\sim$2021
Temperature (T) 13 1.5° 6h Global 1959$\sim$2020
U Component of Wind (U) 13 1.5° 6h Global 1959$\sim$2021
V Component of Wind (V) 13 1.5° 6h Global 1959$\sim$2021
10 Metre U Wind Component (U10M) 1 1.5° 6h Global 1959$\sim$2021
10 Metre V Wind Component (V10M) 1 1.5° 6h Global 1959$\sim$2021
2 Metre Temperature (T2M) 1 1.5° 6h Global 1959$\sim$2021
Mean Sea Level Pressure (MSLP) 1 1.5° 6h Global 1959$\sim$2021
Data used to train $\mathcal{N}_{\theta}$. As shown in Section 3.1, like the atmosphere forcing $\mathrm{A}$, we need another neural network $\mathcal{N}_{\theta}$ to model the ocean surface current dynamics. The data of surface current is obtained from satellie observation, which can be downloaded from https://data.marine.copernicus.eu/product/SEALEVEL_GLO_PHY_L4_MY_008_047. Details of this dataset can be found in Table 5:
Table 5: The data used to train $\mathcal{N}_{\theta}$.
Variable
Name
Spatial
Resolution
Dt
Lat-Lon
Range
Time
Eastward Surface Geostrophic Current Velocity ($\mathrm{U}_{g}$) 0.25° 24h Global 1993$\sim$2021
Northward Surface Geostrophic Current Velocity ($\mathrm{V}_{g}$) 0.25° 24h Global 1993$\sim$2021
Data used to train the hybrid data-driven network. The training of $\dot{\rm S}_{\theta}$ and $\textbf{u}_{\theta}$ involves three atmosphere variables and three ocean varibales, which can be found in Table 6. The SST dataset is selected from GLORYS12V1, which can be downloaded from https://data.marine.copernicus.eu/product/GLOBAL_MULTIYEAR_PHY_001_030.
Table 6: The data used to train $\dot{\rm S}_{\theta}$ and $\textbf{u}_{\theta}$.
Variable
Name
Spatial
Resolution
Dt
Lat-Lon
Range
Time
10 Metre U wind Component (U10M) 0.25° 24h Global 1993$\sim$2021
10 Metre V wind Component (V10M) 0.25° 24h Global 1993$\sim$2021
2 Metre Temperature (T2M) 0.25° 24h Global 1993$\sim$2021
Eastward Surface Geostrophic Current Velocity ($\mathrm{U}_{g}$) 0.25° 24h Global 1993$\sim$2021
Northward Surface Geostrophic Current Velocity ($\mathrm{V}_{g}$) 0.25° 24h Global 1993$\sim$2021
Sea Surface Temperature ($\mathrm{SST}$) 1/12° 24h Global 1993$\sim$2021
B.2 Data preprocessing
Significant magnitude disparities exist among different atmospheric and oceanic variables. To prioritize predictive capability over inter-variable magnitude differences, we implemented standardized preprocessing through normalization. Statistical parameters (mean $\mu$ and standard deviation $\sigma$) were derived from the 1991-2018 training period, with each climate variable being assigned independent scaling coefficients. To be specific, the normalization procedure applied the transformation: $\mathbf{X}^{\prime}=(\mathbf{X}-\mu)/\sigma$, where $\mathbf{X}$ denotes the raw input data.
For observational data, we utlilize the Hadley Centre Integrated Ocean Database (sparse observations) and Global Ocean ODYSSEA L4 Sea Surface Temperature (satellite observations), which can be downloaded from https://climatedataguide.ucar.edu/climate-data/hadiod-met-office-hadley-centre-integrated-ocean-database and https://data.marine.copernicus.eu/product/SST_GLO_PHY_L4_MY_010_044, respectively.
Appendix C Details of Dynamical Core of Ocean-E2E
In this section, we are going to introduce the design of our hybrid physics data-driven model. As shown in Section 3.2, the evolution equation of our hybrid physics data-driven framework can be written as:
$C_{t}=\int_{0}^{t}-({\bf u}(t)+{\bf u}_{\theta}({\bf u}(t),C))\nabla C+\dot{\rm S}_{\theta}({\bf A}(t),C)dt+C_{0}$ (24)
Spatial Discretization. In this study, we utilized two-order center discretization to approximate $\nabla C$. To be sepcific,
$\nabla C(i,j)=({\bf D_{x}}*C(i,j),{\bf D_{y}}*C(i,j))$ (25)
Where and * is convolution operator, ${\bf D_{x}}$ and ${\bf D_{y}}$ are horizontal gradients, which has the form
${\bf D_{x}}=\begin{pmatrix}0&0&0\\ -1/d_{i,j}&0&1/d_{i,j}\\ 0&0&0\end{pmatrix},{\bf D_{y}}=\begin{pmatrix}0&1/s_{i,j}&0\\ 0&0&0\\ 0&-1/s_{i,j}&0\end{pmatrix},$ (26)
$d_{i,j}$ and $s_{i,j}$ are longtitudal and latitdual grid scale at $(i,j)$.
What’s more, unlike atmosphere variables, the ocean MHWs have coastline boundary, which should be treated properly to aviod any numerical instability. In our study, we mitigate this boundary effect through a boundart mask $\mathbf{M}$, where
$\mathbf{M}(i,j)=\begin{cases}1,&\text{if the grid locates on ocean}\\ 0,&\text{else}\end{cases}$ (27)
Furthermore, a scaling factor $\varepsilon_{gm}=0.1$ is employed on ${\bf u}_{gm}$ to avoid the instability of training. The equation 24 then becomes
$C_{t}=\int_{0}^{t}[-({\bf u}(t)+\varepsilon_{gm}{\bf u}_{\theta}({\bf u}(t),C))\nabla C]\mathbf{M}+\dot{\rm S}_{\theta}({\bf A}(t),C)dt+C_{0}$ (28)
Time Integration. As mentioned in Section 3.2, we use second-order Runge-Kutta method to approximate the time integration:
$\tiny C_{(i+1)\Delta t}^{\star}$ $=C_{i\Delta t}+\left[-\left({\bf u}_{i\Delta t}+{\bf u}_{\theta}({\bf u}_{i\Delta t},C_{i\Delta t})\right)\nabla C_{i\Delta t}+\dot{\rm S}_{\theta}(A_{i\Delta t:(i+1)\Delta t},C_{i\Delta t})\right]\Delta t$ (29)
$C_{(i+1)\Delta t}$ $=C_{i\Delta t}+\frac{1}{2}\Bigg{\{}\left[-\left({\bf u}_{i\Delta t}+{\bf u}_{\theta}({\bf u}_{i\Delta t},C_{i\Delta t})\right)\nabla C_{i\Delta t}+\dot{\rm S}_{\theta}(A_{i\Delta t:(i+1)\Delta t},C_{i\Delta t})\right]\Delta t$
$\quad+\left[-\left({\bf u}_{(i+1)\Delta t}+{\bf u}_{\theta}({\bf u}_{(i+1)\Delta t},C_{(i+1)\Delta t}^{\star})\right)\nabla C_{(i+1)\Delta t}^{\star}+\dot{\rm S}_{\theta}(A_{(i+1)\Delta t:(i+2)\Delta t},C_{(i+1)\Delta t}^{\star})\right]\Bigg{\}}$ (30)
In this study, we set $\Delta t=24$ hours and the whole integration time is $t=4$ days.
Appendix D Details of Our Neural Networks
D.1 Multi-scale Convolution
Inspired by [60], this section details the architectures of the neural network modules $\mathbf{u}_{\theta}$ (for subgrid advection) and $\dot{S}_{\theta}$ (for mixing and external source/sink terms) within the Ocean-E2E framework. As shown in Figure 6, both modules leverage the GABlock (Group Attention Block) as their core building unit, designed to efficiently capture multi-scale spatio-temporal features.
Figure 6: Core components of the GABlock. (Top) Core computational flow of the Multilayer Perceptron (Mlp) module, including a 1x1 convolution, a depthwise convolution, GELU activation, and Dropout. (Bottom) Core computational flow of the Spatial Attention (SA) module, illustrating the generation of feature $F_{x}$ and gating signal $G_{x}$ through an initial 1x1 convolution, GELU activation, and a depthwise convolution. The gating signal $G_{x}$ is then passed through a Sigmoid activation and element-wise multiplied with $F_{x}$. The result is processed by a final 1x1 convolution and integrated with the input via a residual connection. These two components collectively form the key information processing units within the GABlock.
[…]
$\mathbf{F}_{g}=\text{Conv}_{1\times 1}\left(\text{DWDConv}_{\text{spatial}}\left(\text{DWConv}_{0}\left(\sigma_{\text{act}}\left(\mathcal{P}_{1}\left(\mathbf{X}_{\text{norm1}}\right)\right)\right)\right)\right)$ (31)
$\mathbf{F}_{x},\mathbf{G}_{x}=\text{split}\left(\mathbf{F}_{g},\text{dim}=1\right)$ (32)
$\text{SGU}_{\text{out}}=\sigma_{\text{sigmoid}}\left(\mathbf{G}_{x}\right)\odot\mathbf{F}_{x}$ (33)
The result from the SA module, $\mathcal{P}_{2}(\text{SGU}_{\text{out}})$, is integrated back into the input via a residual connection, scaled by a learnable parameter $\lambda_{1}$: $\mathbf{X}_{\text{attn}}=\mathbf{X}+\lambda_{1}\cdot\text{DropPath}(\mathcal{P}_{2}(\text{SGU}_{\text{out}}))$. Subsequently, $\mathbf{X}_{\text{attn}}$ undergoes another normalization to $\mathbf{X}_{\text{norm2}}=\text{BN}(\mathbf{X}_{\text{attn}})$ and is processed by a Multilayer Perceptron (Mlp) module. The Mlp typically consists of two 1x1 convolutions sandwiching a depth-wise convolution, along with activations and dropout. Its output, $\text{Mlp}(\mathbf{X}_{\text{norm2}})$, is also added residually with a learnable scale $\lambda_{2}$: $\mathbf{X}_{\text{out}}=\mathbf{X}_{\text{attn}}+\lambda_{2}\cdot\text{DropPath}(\text{Mlp}(\mathbf{X}_{\text{norm2}}))$.
The module for approximating the bolus velocity, $\mathbf{u}_{\theta}$, adopts a U-Net-like encoder-decoder architecture. It takes the concatenated ocean surface current $\mathbf{u}\in\mathbb{R}^{B\times 2\times H\times W}$ and SSTA $\mathbf{C}\in\mathbb{R}^{B\times 1\times H\times W}$ (forming $\mathbf{X}_{u}\in\mathbb{R}^{B\times 3\times H\times W}$) as input. The encoder path, composed of $L_{e}$ stages, utilizes downsampling convolutions (e.g., ConvSC) and GABlocks to extract multi-scale features, denoted as $\mathcal{H}_{l}$ for the $l$-th stage. Deeper encoder layers, particularly the bottleneck which processes $\mathcal{H}_{L_{e}}$ with stacked GABlocks, are crucial for capturing non-local spatial dependencies inherent in subgrid effects. The decoder path, with $L_{d}$ stages, symmetrically reconstructs spatial resolution using upsampling convolutions and GABlocks, integrating features from corresponding encoder stages $\mathcal{H}_{L_{e}-l+1}$ via skip connections. A final convolutional layer then maps the decoder’s output to the bolus velocity field $\mathbf{u}_{\theta}\in\mathbb{R}^{B\times 2\times H\times W}$.
The module for parameterizing mixing and external source/sink terms, $\dot{S}_{\theta}$, primarily consists of a stack of $N_{s}$ GABlocks. Its input is the concatenation of atmospheric forcing variables $\mathbf{A}\in\mathbb{R}^{B\times C_{A}\times H\times W}$ and SSTA $\mathbf{C}\in\mathbb{R}^{B\times 1\times H\times W}$ (forming $\mathbf{X}_{s}\in\mathbb{R}^{B\times(C_{A}+1)\times H\times W}$). An initial convolution, InitialConv, first adjusts $\mathbf{X}_{s}$ to a hidden channel dimension $C_{\text{hid}}$. This is followed by $N_{s}$ serially stacked GABlockl ($l=1,\dots,N_{s}$). Let $\mathcal{H}_{l}$ denote the output of the $l$-th GABlock. Finally, an OutputConv layer maps the features from the last GABlock, $\mathcal{H}_{N_{s}}$, to a single-channel output representing the SSTA tendency $\dot{S}_{\theta}\in\mathbb{R}^{B\times 1\times H\times W}$.
These architectures, centered around the versatile GABlock, enable $\mathbf{u}_{\theta}$ and $\dot{S}_{\theta}$ to effectively learn and represent the complex physical processes governing SSTA evolution directly from observational and reanalysis data.
D.2 Neural Data Assimilation Network
Inspired by  [38], we apply an edge-reparameterization and attention-guided network to project the observation to analysis data, which achieves efficent assimilation for real-time application. It primarily consists of four parts, i.e., a convolutional layer (ConvL), a Kirsch-guided reparameterization module (KRM), a channel attention module (CAM), and a spatial attention module (SAM).
D.2.1 Convolutional Layer
The convolutional layer (ConvL) can defined as:
$\mathrm{ConvL}(x_{in})=\mathcal{PR}(\mathcal{LN}(Conv(x_{in}))),$ (34)
where $x_{in}\in\mathbb{R}^{C\times H\times W}$ is the input of $\mathrm{ConvL}$. $\mathrm{Conv}$, $\mathcal{LN}$, and $\mathcal{PR}$ represent the convolution operation, the layer normalization, and parametric rectified linear unit (PReLU), respectively.
D.2.2 Attention Mechanism
The attention mechanism can realize the efficient allocation of information processing resources and can give more attention to key area of observatiuon while temporarily ignoring the missing locations. Therefore, the attention mechanism is used to focus on important information with high weights, ignore irrelevant information with low weights, and continuously adjust the weights during the network learning process. Therefore, a single model can extract more valuable feature information in different imaging environments. In this work, both channel attention and spatial attention are jointly exploited to further improve the assimilation performance.
Channel Attention Module. The channel attention mechanism is thus exploited to reconstruct the relationship between target analysis features and the input of different observation fields. The channel attention module ($\mathrm{CAM}$) can be defined as:
[…]
where $x_{in}^{c}$ is the input of $\operatorname{CAM}$, $\operatorname{MLP}$ denotes the multilayer perceptron, and $\sigma$ is the Sigmoid nonlinear activation function.
Spatial Attention Module. In data assimilation, the spatial attention module (SAM) is used to concentrate on crucial regions, which can be defined as:
$\operatorname{SAM}$ $=\sigma\left(\mathrm{Conv}^{7\times 7}([\mathrm{Avg}(x_{in}^{s});\mathrm{Max}(x_{in}^{s})])\right),$ (36)
[…]
D.2.3 Kirsch-guided Reparameterization Module
To facilitate efficient assimilation in real-world MHWs forecasting, we employ the Kirsch-guided reparameterization module (KRM) with shared parameters, as proposed by Liu et al. [38]. The KRM is composed of three principal elements: (1) a standard convolutional layer that captures local spatial features; (2) an expand–squeeze convolutional layer that adaptively adjusts the receptive field; and (3) a family of eight directional edge-detection operators that guide the learning and inference of the convolutional kernels. Specifically, the standard convolutional and expand–squeeze operations can be formulated as follows:
$\begin{gathered}F_{n}=W_{n}*x_{in}^{k}+B_{n},\end{gathered}$ (37)
$\begin{gathered}F_{es}=W_{s}*\left(W_{e}*x_{in}^{k}+B_{e}\right)+B_{s},\end{gathered}$ (38)
where $x_{in}^{k}$ denotes the input of KRM. $W_{n}$, $W_{s}$, $W_{e}$, $B_{n}$, $B_{s}$, $B_{e}$ are the weights and bias of convolution operation. $F_{n}$ represents the output of the normal convolutional layer. $F_{es}$ denotes the output of the expanding-and-squeezing convolutional layer. We first reparameterize the Eqs. (37) and (38). Subsequently, we merge them into a single normal convolution with parameters $W_{es}$ and $B_{es}$:
$W_{es}=\mathrm{perm}\left(W_{e}\right)*W_{s},$ (39)
$B_{es}=W_{s}*\mathrm{rep}\left(B_{e}\right)+B_{s},$ (40)
where $\operatorname{perm}(\cdot)$ denotes the permute operation which exchanges the $1$st and $2$nd dimensions of a tensor, $\operatorname{rep}(\cdot)$ denotes the spatial broadcasting operation. And the predefined eight-direction Kirsch edge filters $K_{i}$ are incorporated into the reparameterization module. To memorize the edge features, the input feature $x_{in}^{k}$ will first be processed by $C\times C\times 1\times 1$ convolution and then use a custom Kirsch filter to extract the feature map gradients in eight different directions. Therefore, the edge information in eight directions can be expressed as follows:
$F_{K}^{i}$ $=\left(S_{K}^{i}\odot K_{i}\right)\otimes\left(W_{i}*x_{in}^{k}+B_{i}\right)+B_{K_{i}}$ (41)
$=W_{K}^{i}\otimes\left(W_{i}*x_{in}^{k}+B_{i}\right)+B_{K_{i}},$
where $W_{i}$ and $B_{i}$ are the weights and bias of $1\times 1$ convolution for branches in eight directions, $S_{K}^{i}$ and $B_{K_{i}}$ are the scaling parameters and bias with the shape of $C\times 1\times 1\times 1$, $\odot$ indicates the channel-wise broadcasting multiplication. The combined edge information, extracted by the scaled Kirsch filters, is given by:
$F_{\mathrm{K}}=\sum_{i=1}^{8}F_{K}^{i}.$ (42)
Therefore, the final weights and bias after reparameterization can be expressed as:
$W_{\mathrm{rep}}$ $=W_{n}+W_{es}+\sum_{i=1}^{8}\left(\mathrm{perm}\left(W_{i}\right)*W_{\mathrm{K}}^{i}\right),$ (43)
$B_{\mathrm{rep}}$ $=B_{n}+B_{es}+\sum_{i=1}^{8}\left(\mathrm{perm}\left(B_{i}\right)*B_{\mathrm{K}_{i}}\right).$ (44)
Finally, the output can be obtained using a single normal convolution in the inference stage:
$F=W_{\mathrm{rep}}*x_{in}^{k}+B_{\mathrm{rep}}.$ (45)
D.2.4 The Assimilation Netwrok
We fist define the basic block:
$res_{1}=\mathcal{PR}(\mathcal{G}(Conv(x_{in}^{b}))),$ (46)
$res_{2}=\mathcal{PR}(\mathcal{G}(Conv(res_{1}))),$ (47)
$res_{attention}=SAM(CAM(res2)),$ (48)
$res_{3}=\mathcal{PR}(\mathcal{G}(Conv(res_{attention}))+x_{in}^{b}),$ (49)
$res_{4}=\mathcal{PR}(\mathcal{G}(Conv(res_{3}))),$ (50)
where, $x_{in}^{b}$ is the input of basic block, $\mathcal{G}$ represents GroupNorm. In summary, the basic block can be defined as:
$res_{4}=Block(x_{in}^{b}).$ (51)
Therefore, for the obsevation data $x_{o}$, it can be processed to analysis data as follow:
$y_{a}^{1}=Block(KRM(x_{o})+x_{o}),$ (52)
$y_{a}^{2}=Block(KRM(y_{a}^{1})+y_{a}^{1}),$ (53)
$y_{a}^{3}=Block(KRM(y_{a}^{2})+y_{a}^{2}),$ (54)
$y_{a}^{4}=Block(KRM(y_{a}^{3})+y_{a}^{3}),$ (55)
$y_{a}^{5}=Block(KRM(y_{a}^{4})+y_{a}^{4}),$ (56)
$y_{a}^{out}=\mathcal{C}(y_{a}^{5}),$ (57)
where, $\mathcal{C}$ denotes the final convolution operation, which maps the latent information to the analysis data.
Appendix E Experiments Details
E.1 Evaluation Metric
Definition of MHWs. The [25] definition of MHW categories is applied to detect MHWs from original SST field. More precisely, any interval in which the SST anomalies surpass the 90th percentile for a minimum of five days in a row is classified as an MHW event. Subsequently, the MHWs are divided into three severity categories: moderate, strong, and severe/extreme. These categories are determined by the 90th, 92.5th, and 95th percentiles of the SST, respectively [20]. Note that our results are unaffected by different definitions of the MHW category.
Metrics. In this study, to evaluate the performance of our model, we use three metrics: Root Mean Square Error (RMSE), Critical Success Index (CSI) and Bias (BIAS). RMSE represents the overall performance of our SST anomaly forecast, which is calculated as:
${\rm RMSE}(t)=\sqrt{\frac{1}{H\times W}\sum_{i,j}\omega_{i,j}(T_{i,j}(t)-\tilde{T}_{i,j}(t))^{2}}$ (58)
Where $T_{i,j}(t)$ and $\tilde{T}_{i,j}(t)$ is the ground truth and prediction of SST anomaly at lead time $t$, $\omega_{i,j}$ is the weighting coefficient and the subscript $(i,j)$ indicates the data at grid point $(i,j)$ while $H$ and $W$ follow the same definition in Section 2. CSI is the metric that evaluate the forecast ability of extreme events and can be expressed as:
${\rm CSI}(t)=\frac{{\rm TP}}{{\rm TP}+{\rm FP}+{\rm FN}}$ (59)
Where TP (True Positive) denotes the number of cases that a MHW event is accurately predicted. FP (False Positive), FN (False Negative), and TN (True Negative) follow a similar definition.
BIAS is used to evaluate the systematic difference between the analysis field and the ground truth:
${\rm BIAS}(t)=\frac{1}{H\times W}\sum_{i,j}\omega_{i,j}(T_{i,j}(t)-\tilde{T}_{i,j}(t))$ (60)
E.2 Model Training
All baseline models and Ocean-E2E were trained under identical experimental configurations to ensure fair comparison. The training process employed 100 epochs with an initial learning rate of $1\times 10^{-3}$, utilizing a step-wise learning rate scheduler to dynamically adjust training parameters until convergence was achieved. For objective evaluation, model checkpoints demonstrating optimal performance on the validation set were selected as final candidates for comparative analysis.
Appendix F Related Work and Limitations
F.1 Numerical Model based Ocean Forecasting
Traditional MHWs forecasting predominantly relies on numerical models solving oceanic primitive equations. These approaches are typically categorized into two types: (1) seasonal forecasting, which predicts large-scale MHW characteristics including onset timing, duration, and intensity evolution [30, 8], and (2) sub-seasonal forecasting, which resolves finer spatiotemporal variations at the cost of reduced prediction horizons [2, 63]. While physics-based models demonstrate reasonable forecast skill, they face two principal limitations: prohibitive computational costs scaling with model resolution [62, 58], and systematic underestimation of extreme MHW magnitudes [30].
F.2 Deep Learning based Ocean Forecasting
Machine learning (ML) techniques have emerged as promising alternatives for oceanic modeling, offering orders-of-magnitude speedup over conventional methods [37, 22]. Recent efforts have developed spatio-temporal architectures including NMO [59], neXtSIM [16], and OceanVP [50] for regional ocean modeling, alongside specialized models for specific oceanic phenomena such as ENSO [21, 12], MJO [31, 14, 51], and MHWs [49, 30, 52]. Building on advances in atmospheric foundation models [44, 4], the community has recently proposed global ocean foundation models like AI-GOMS [62], Xihe [58], and WenHai [13].
F.3 Deep Learning based Data Assimilation
Deep learning (DL) has shown transformative potential in weather and climate data assimilation (DA), with progress evolving along three directions: (1) Conceptual validation using idealized systems like Lorenz models and shallow water equations [7, 1, 15, 33]; (2) Component replacement within conventional DA frameworks, including physical parameterization [23], observation operators [35, 54], and optimization cost functions [41, 61]; (3) End-to-end DA system replacement, exemplified by FengWu-Adas and DiffDA that implement complete DL-based assimilation pipelines using ERA5 reanalysis data [11, 27].
Appendix G Additional Results
G.1 Additional Comparison with S2S System
In this section, we are going to show additional comparison results with S2S system. To be specific, we take the S2S analysis field as the groundtruth, and use it to initialize our Ocean-E2E. The atmospheric boundary condition is acquired through one of the state-of-the-art AI-driven weather forecast model: OneForecast [17]. As shown in Table 7, our model still performs better than S2S.
Table 7: Additional comparison with S2S. The average results for global SSTA of RMSE are recorded. A small RMSE ($\downarrow$) indicates better performance. The best results are in bold, and the second best are with underline.
Model Metric (RMSE)
Time Interval
14-day 21-day 28-day 35-day
S2S (2020) [44] 0.5249 0.5994 0.6618 0.7086
Ocean-E2E (2020) [39] 0.4321 0.5242 0.5891 0.6304
Ocean-E2E (Promotion) 17.8 % 12.5 % 11.0 % 11.0 %
S2S (2021) [44] 0.5175 0.5703 0.6121 0.6448
Ocean-E2E (2021) [39] 0.4350 0.5167 0.5626 0.6022
Ocean-E2E (Promotion) 15.9 % 9.4 % 8.1 % 6.6 %
G.2 Visualization of Global MHWs Simulation
In this section, we visualize the global MHWs forecast result of our Ocean-E2E and S2S, as shown in Figure 7-10. We plot the global SSTA field, where darkred parts indicate strong MHWs.
G.3 Visualization of Global MHWs Forecast
In this section, we visualize the global MHWs forecast result of our Ocean-E2E and S2S, as shown in Figure 11-14. We plot the global SSTA field, where darkred parts indicate strong MHWs.
G.4 Visualization of Regional High Resolution Simulation
In this section, we visualize the regional (North Pacific) simulation of $C_{t}-C_{0}$, as shown in Figure 15-18. We plot the regional SSTA field, where darkred parts indicate strong intensifications of the MHWs.
Figure 7: Global Simulation of MHWs Initialized on January 13, 2020
Figure 8: Global Simulation of MHWs Initialized on May 12, 2020
Figure 9: Global Simulation of MHWs Initialized on July 11, 2020
Figure 10: Global Simulation of MHWs Initialized on October 9, 2020
Figure 11: Global Forecast of MHWs Initialized on January 13, 2020
Figure 12: Global Forecast of MHWs Initialized on May 12, 2020
Figure 13: Global Forecast of MHWs Initialized on July 11, 2020
Figure 14: Global Forecast of MHWs Initialized on October 9, 2020
Figure 15: Regional Simulation of North Pacific Initialized on January 13, 2020
Figure 16: Regional Simulation of North Pacific Initialized on May 12, 2020
Figure 17: Regional Simulation of North Pacific Initialized on July 11, 2020
Figure 18: Regional Simulation of North Pacific Initialized on October 9, 2020